A constant force of 40 N is applied by a merchandiser on 6 seater dining table along a smooth horizontal plane. The dining table starts from rest and observed to move 100 meters in 5 seconds. . Find a) acceleration of the dining table b) mass of the dining table.


1
Expert's answer
2021-04-01T18:41:54-0400

Solution:


assuming zero friction.


d = "\frac{ at²}{2}"

100 = "\frac{a5²}{2}"

a  = "\frac{200}{25}" = 8 "(\frac{m}{s^2})"


F = ma


m = "\dfrac{F}{a}=\dfrac{40}{8}=5N"